### 13 May 2026 v1

# Hierarchical Transformer Preconditioning for Interactive Physics Simulation

Abstract.
Neural preconditioners for real-time physics simulation offer promising data-driven priors, but they often fail to capture long-range couplings efficiently because they inherit local message passing or sparse-operator access patterns. We introduce the Hierarchical Transformer Preconditioner, a neural preconditioner anchored to a weak-admissibility $\mathcal{H}$-matrix partition. The partition provides a multiscale structural prior (dense diagonal leaves plus coarsening off-diagonal tiles) that enables full-graph approximate-inverse computation with $O(N)$ scaling at fixed block sizes. The network models the inverse through low-rank far-field factors and uses highway connections (axial buffers plus a global summary token) to propagate context across transformer depth. At each PCG iteration, preconditioner application reduces to batched dense GEMMs with regular memory access.
The key training contribution is a cosine-Hutchinson probe objective that learns the action of MAMA on convergence-critical spectral subspaces, optimizing angular alignment of $MA\mathbf{z}$ with $\mathbf{z}$ rather than forcing eigenvalue clusters to a prescribed location. This removes unnecessary spectral-placement constraints from SAI-style objectives and improves conditioning on irregular spectra. Because both inference and apply are dense, dependency-free tensor programs, the full solve loop is captured as a single CUDA Graph.
On stiff multiphase Poisson systems (up to 100:1100\!:\!1 density contrast, N=1,024N\!=\!1{,}02416,38416{,}384), the solver runs from $\sim\!143$ to $\sim\!21$ fps. At N=8,192N\!=\!8{,}192, it reaches $17.9$ ms/frame, with $2.2\times$ speedup over GPU Jacobi, $\sim\!28\times$ over GPU IC/DILU (AMGX multicolor_dilu), and $2.7\times$ over neural SPAI retrained per scale on the same benchmark.
preconditioning, iterative solvers, hierarchical matrices, transformers, neural operators, physics simulation, real-time graphics
copyright: nonejournalyear: 2026doi: 10.1145/nnnnnnn.nnnnnnnjournal: TOGjournalvolume: 45journalnumber: 6article: 1publicationmonth: 12ccs: Computing methodologies Physical simulationccs: Computing methodologies Neural networks
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1. Teaser. Hierarchical neural preconditioning for interactive multiphase Poisson solves: (a) weak-admissibility $\mathcal{H}$-matrix prior, (b) learned preconditioner $M\!\approx\!A^{-1}$, (c) multiscale residual transport induced by $M$, and (d) 3D multiphase fluid simulation application context.
1. Introduction
Real-time simulation of fluids, soft bodies, and coupled multiphysics often reduces every time step to a large sparse SPD solve attacked with preconditioned conjugate gradient (PCG) (Saad, 2003). The preconditioner $M\!\approx\!A^{-1}$ is the dominant design variable. Classical choices — Jacobi, incomplete Cholesky/LU (IC/ILU), sparse approximate inverses (SPAI), and algebraic multigrid (AMG) (Lin and Moré, 1999; Kolotilina and Yeremin, 1993; Benzi et al., 1996; Ruge and Stüben, 1987) — excel when the matrix is reused across many solves so a heavy setup amortizes. In an interactive setting AA changes every frame, so any preconditioner whose setup approaches the per-frame budget cannot amortize; IC and AMG suffer worst (Naumov et al., 2015; Liu et al., 2016; Yamazaki et al., 2020).
Machine learning offers an alternative: train a neural preconditioner once and amortize the cost across frames. Graph neural networks have been the dominant choice (Li et al., 2023; Häusner et al., 2023; Chen, 2024; Yang et al., 2025) because a sparse matrix has a natural graph interpretation, but the choice forces two limitations that become acute in real-time simulation. (i) GNN preconditioners typically inherit the sparsity of AA, so non-adjacent interactions are either dropped or only reached through repeated graph convolutions that oversmooth local detail (Wu et al., 2019; Trifonov et al., 2025). (ii) The dominant operations at apply time — triangular solves on a learned IC, or two SpMVs on a learned SPAI — are bandwidth-bound and either serialize on data hazards or scatter through irregular memory.
From PCG geometry to hierarchical structure.
The ideal preconditioner is $A^{-1}$, which is dense even when AA is sparse, so the question is what structure makes $A^{-1}$ cheap to store and apply. In modern simulation codes the degrees of freedom are already laid out along a space-filling curve (Morton, Hilbert) or a bandwidth-reducing permutation, because the surrounding pipeline (spatial hashing, BVH, neighbor search) wants that ordering (Karras, 2012; Teschner et al., 2003; Ihmsen et al., 2011). Under such orderings the nonzeros of AA cluster near the diagonal and the corresponding off-diagonal blocks of $A^{-1}$ describe long-range interactions whose numerical rank decays with separation. Hierarchical matrices ($\mathcal{H}$-matrices) (Hackbusch, 1999; Hackbusch and Khoromskij, 2002; Börm et al., 2003) exploit exactly this — a recursive block partition with dense diagonal blocks and low-rank off-diagonal blocks — and the fast multipole method (Greengard and Rokhlin, 1987) factors the same prior for translation-invariant kernels. An $\mathcal{H}$-matrix preconditioner is architecturally an FMM-shaped operator on the assembled system, and that is the prior we hand the network.
Approach.
We anchor the preconditioner to a weak-admissibility $\mathcal{H}$-matrix partition (Hackbusch and Khoromskij, 2002), computed analytically from the leaf-index geometry and reused for every frame: K=N/LK\!=\!N/L dense diagonal blocks of size $L\!\times\!L$ plus off-diagonal tiles spanning $S\!\times\!S$ leaves with $S\!=\!2^{j}$. A two-stream transformer operates on this layout. A diagonal stream emits one factor per leaf at full rank; an off-diagonal stream emits one factor per tile at a constant coarse token count $L_{s}\!\ll\!L$ regardless of $S$, so the rank fraction $L_{s}/(SL)$ shrinks automatically with separation — exactly the bias $\mathcal{H}$-matrix compression calls for. To let the network route information between blocks without breaking the $\Theta(N)$ budget, we add per-layer highway buffers (axial row/column bands plus a single global token, §4.3).
A loss with no preferred eigenvalue location.
We train self-supervised with a cosine-similarity Hutchinson probe loss on the preconditioned image $MA\mathbf{z}$. The conceptual move over the Frobenius/SAI loss used by (Yang et al., 2025) is to replace a distance between vectors ($\|\tfrac{1}{\|A\|}AM\mathbf{z}-\mathbf{z}\|_{2}$) with a distance between subspaces (the angle between the lines spanned by $\mathbf{z}$ and $MA\mathbf{z}$). This is invariant under positive rescaling of $M$, exactly matching PCG, which cares about the direction MAMA sends each probe and not the magnitude. The network is then free to cluster the spectrum of MAMA wherever angular alignment is easiest, with no implicit demand that the cluster live near any particular eigenvalue (Prop. 6.1; spectral evidence in Fig. 2). Empirically, this is not a minor objective tweak: with architecture and training split fixed, the loss swap alone accounts for a major quality gap in Table 6.
One CUDA Graph for the whole iteration.
The choices above together remove the data dependencies that normally fragment a CG iteration into separately dispatched kernels: the partition is fixed, every tile shape is the same, the apply is batched GEMMs over preallocated tensors with no triangular solves and no allocations. The whole PCG inner loop — preconditioner apply, SpMV, SAXPYs, dot reductions — records into one CUDA Graph that replays per iteration with zero CPU dispatch overhead. In our measurements this is a substantial share of the gap to classical GPU preconditioners.
Contributions.
In short: [c1] a neural preconditioner that bakes the $\mathcal{H}$-matrix block partition in as a structural prior with a fixed coarse rank on every off-diagonal tile, keeping total work at $\Theta(N)$ (intro paragraph c1, §4.2); [c2] highway buffers that route information between blocks without growing the budget (§4.3); [c3] a cosine-similarity Hutchinson loss that minimizes a distance between subspaces rather than between vectors, exactly matching PCG’s scale-invariance and freeing the network to put the spectrum of MAMA wherever convergence is easiest (intro paragraph c3, Prop. 6.1 in §6.1); [c4] an allocation-free apply with no triangular solves, no data-dependent branching, no per-kernel CPU dispatch — the whole PCG inner loop captured as one CUDA Graph (intro paragraph c4, §5). Table 1 summarizes the per-iteration access pattern and throughput limit of each preconditioner family at the one step where every fixed-pattern alternative loses time on modern GPUs.
Table 1. Per-iteration preconditioner application: arithmetic cost, dominant memory-access pattern, GPU throughput limit, and whether the apply admits single-CUDA-Graph capture of the PCG inner loop. $\mathrm{nnz}$ is the nonzero count of the relevant factor. $L_{s}$” is our fixed coarse-token count per off-diagonal tile.
Method FLOPs Memory access GPU bottleneck Graph
Jacobi $O(N)$ stride-1 read memory bandwidth
IC (tri. solve) $O(\mathrm{nnz})$ wavefront-dep. SpTRSV latency / data hazards $\times$
Neural SPAI (Yang et al., 2025) $O(\mathrm{nnz})$ random gather (SpMV) memory bandwidth
AMG (V-cycle) $O(N\log N)$ multilevel, irregular level synchronization $\times$
Ours $O(NL_{s})$ batched GEMM, stride-1 compute (Tensor Cores)
2. Related Work
Classical preconditioners.
Jacobi, incomplete Cholesky/LU, sparse approximate inverses (SPAI), and algebraic multigrid (Saad, 2003; Lin and Moré, 1999; Kolotilina and Yeremin, 1993; Benzi et al., 1996; Ruge and Stüben, 1987; Naumov et al., 2015) are mature and remain default choices when the system is reused across many solves so a careful setup amortizes. The known GPU costs of triangular solves (Liu et al., 2016; Yamazaki et al., 2020) and of V-cycle communication (Naumov et al., 2015) bite hardest in the regime we target (AA changing every frame at interactive rates), which motivates the amortized neural alternative pursued here.
Learned PDE solvers and preconditioners.
Neural operators (Fourier (Li et al., 2021), DeepONet (Lu et al., 2021), graph-based (Li et al., 2020b, a)) and physics-informed networks (Raissi et al., 2019) learn function-space mappings between coefficient and solution fields, and have been used both as surrogates and as inner preconditioners for flexible Krylov solvers (Rudikov et al., 2024); we operate one rung lower, on the assembled algebraic system. A productive line of work trains GNNs to emit a learned IC factor (Li et al., 2023; Häusner et al., 2023; Trifonov et al., 2024) or a sparse approximate inverse in factored $GG^{\top}$ form using a scale-invariant Frobenius (SAI) loss (Yang et al., 2025); we compare to the latter directly in Table 4. Chen (Chen, 2024) reports strong performance across SuiteSparse problems where classical IC and AMG struggle, and Trifonov et al. (Trifonov et al., 2025) argue that message-passing GNNs cannot approximate the non-local elimination structure of sparse triangular factors — a motivation for architectures (like ours) that route global information through explicit channels.
Hierarchical neural and matrix machinery.
[…]
3. Background
CG converges through the polynomial bound
(1) $\|\mathbf{e}_{k}\|_{A}\;\leq\;\|\mathbf{e}_{0}\|_{A}\,\min_{p\in\mathcal{P}_{k},\,p(0)=1}\;\max_{\lambda\in\sigma(A)}|p(\lambda)|,$
so CG only sees AA through how small a polynomial fixed to 11 at the origin can be made on $\sigma(A)$. The practical consequence we lean on is that when $\sigma(MA)$ collapses to $c$ tight clusters away from zero, CG converges in $\sim\!c$ iterations regardless of where on the real line those clusters sit. The preconditioning objective is therefore “cluster $\sigma(MA)$,” not “make MAMA close to II,” with cluster location essentially free — the property our cosine loss exploits.
An $\mathcal{H}$-matrix (Hackbusch, 1999; Hackbusch and Khoromskij, 2002; Börm et al., 2003) represents a dense matrix as a recursive block partition: dense diagonal blocks plus off-diagonal blocks stored as low-rank factors $UV^{\top}$, with admissible rank shrinking off the diagonal because the underlying Green’s function is smooth there. We use the weak-admissibility variant (Hackbusch and Khoromskij, 2002; Hartland et al., 2023), whose fewer, larger off-diagonal tiles batch well on a GPU; classical $\mathcal{H}$-matrix arithmetic costs $O(N\log N)$, our learned realization runs at $\Theta(N)$ thanks to the fixed coarse-token count in c1.
The training loss uses Hutchinson-style probes $\mathbf{z}$ with $\mathbb{E}[\mathbf{z}\mathbf{z}^{\top}]\!=\!I$ (Hutchinson, 1989), but not as a trace estimator. The property we use is spectral whiteness: isotropy gives $\mathbb{E}[(\mathbf{u}_{i}^{\top}\mathbf{z})^{2}]\!=\!1$ for every eigenvector $\mathbf{u}_{i}$ of MAMA, so a scalar built from $(\mathbf{z},MA\mathbf{z})$ weighs every eigenmode equally rather than favoring smooth or oscillatory ones — the property that makes the cosine loss of c3 a faithful global indicator of preconditioner quality.
4. Method
Notation and pipeline.
$A\!\in\!\mathbb{R}^{N\times N}$ is the assembled sparse SPD system, indexed along a space-filling-curve or bandwidth-reducing ordering. The NN indices are partitioned into K=N/LK\!=\!N/L contiguous leaves of size $L$, inducing an $\mathcal{H}$-matrix partition of $K$ diagonal blocks and $M_{\mathcal{H}}$ unique weakly-admissible off-diagonal tiles (tile $m$ spans $S_{m}\!\times\!S_{m}$ leaves). The network is a four-stage pipeline: a physics-aware encoder, a diagonal attention stack (full per-node resolution $L$), an off-diagonal attention stack (coarse token count $L_{s}\!\ll\!L$, the same for every tile), and linear decoder heads. The whole network dispatches exactly two attention kernels regardless of NN — one batched over leaves, one batched over tiles — which, together with the static partition, is what later lets the entire PCG inner loop run inside a single CUDA Graph. We use $L\!=\!128,L_{s}\!=\!32,d\!=\!128,n_{l}\!=\!3$ in every reported result.
4.1. Encoder
A two-layer MLP lifts whichever per-node features the simulator already exposes (position, velocity, density, pressure, material parameters) plus the geometric and coupling features $(\Delta\mathbf{x}_{ij},A_{ij})$ for every nonzero of AA to width-$d$ node embeddings; $n_{\mathrm{gcn}}\!=\!2$ residual graph-convolutional layers (Kipf and Welling, 2017) mix in neighborhood information using AA as the message-passing weight. This is the only stage that observes individual graph edges; everything downstream operates on the block layout of the partition, decoupling cost from $\mathrm{nnz}(A)$. Because $M$ is neural rather than algebraic, the network can look at $(\rho_{i},\Delta\mathbf{x}_{ij})$ directly — physical structure a fixed analytical preconditioner cannot see (full feature list in Supplementary LABEL:SX-supp:architecture).
4.2. Diagonal and off-diagonal attention stacks
Diagonal stack. For each of the $K$ leaves, $n_{l}$ transformer layers (Liu et al., 2021; Vaswani et al., 2017) with within-leaf (Swin-style) attention and edge-bias logits emit a dense factor $F_{k}\!\in\!\mathbb{R}^{L\times L}$. The corresponding diagonal block of $M$ is the PSD outer product $F_{k}F_{k}^{\top}$.
Off-diagonal stack. For each off-diagonal tile $m$ of size $S_{m}L\!\times\!S_{m}L$, node embeddings on its row- and column-strips are pooled to a fixed coarse token count $L_{s}\!=\!L/p_{\mathrm{off}}\!\ll\!L$ regardless of $S_{m}$. $n_{l}$ transformer layers over these $L_{s}$ tokens emit a factor pair $U_{m},V_{m}\!\in\!\mathbb{R}^{L_{s}\times L_{s}}$; the tile’s coarse representation is $B_{m}\!=\!U_{m}V_{m}^{\top}$. Because the token count is the same on every tile, the implied rank fraction $L_{s}/(S_{m}L)$ shrinks automatically with separation — exactly the $\mathcal{H}$-matrix prior that distant blocks need less rank. A rank-fraction audit on real frames confirms the architecture-provided $L_{s}/(SL)$ stays above what truncated SVD needs at $\varepsilon\!\in\!\{10^{-3},10^{-6},10^{-9}\}$ in every distance class (Fig. 3; assembled-$M$ visualization in Supplementary, Fig. LABEL:SX-supp-fig:matrices). With $S_{m}$ doubling geometrically, $M_{\mathcal{H}}\!=\!O(K)$ unique off-diagonal tiles, and a tile-batched attention of fixed shape, the total work over the off-diagonal stack stays linear in NN.
4.3. Highway connections
Within-block attention is local by construction. To let the network route information across the matrix without giving up the $\Theta(N)$ budget, after every attention sublayer we scatter-add block-token embeddings into three buffers per transformer layer: a row-band axial buffer $\mathbf{r}_{\mathrm{hw}}$, a column-band axial buffer $\mathbf{c}_{\mathrm{hw}}$, and a single global summary token $\mathbf{g}_{\mathrm{hw}}$. Each token then concatenates its row, column, and global context with its own embedding before the FFN sublayer. The four channels per layer (2D intra-block, 1D row, 1D column, 0D global) cost $O(Nd)$ scatter-gather and preserve $\Theta(N)$ scaling. We ablate the highways in §7.3 and illustrate the per-layer connectivity in Supplementary, Fig. LABEL:SX-supp-fig:partition-and-highwaysb.
$4.$ Decoder, geometric picture, and complexity
Three linear decoder heads project tokens to the final factors: a leaf head emits $F_{k}\!\in\!\mathbb{R}^{L\times L}$ (diagonal block $F_{k}F_{k}^{\top}$, PSD by construction); two off-diagonal heads emit $U_{m},V_{m}\!\in\!\mathbb{R}^{L_{s}\times L_{s}}$; two node heads emit per-leaf bridges $\tilde{U}_{k},\tilde{V}_{k}\!\in\!\mathbb{R}^{L\times L_{s}}$ between node resolution and coarse tile resolution. A small learned per-node Jacobi gate $\lambda_{i}$ adds a diagonal correction $\mathrm{diag}(\boldsymbol{\lambda})\mathrm{diag}(A)^{-1}$ that absorbs gradient early in training and decays toward zero as the structured branches take over (Supplementary LABEL:SX-supp:architecture). In reported runs we do not explicitly enforce strict SPD and did not observe instability; if a strict guarantee is required, one can add a tiny positive diagonal shift (e.g., softplus-parameterized, as in prior learned preconditioners (Yang et al., 2025)). All factors pack into a single tensor of width $P\!=\!KL^{2}+M_{\mathcal{H}}L_{s}^{2}+2NL_{s}+N$ that the apply consumes without ever materializing $M$.
Complexity and measured breakdown.
Encoder cost is $O(Nd^{2})$, diagonal stack $O(n_{l}NLd)$, off-diagonal stack $O(n_{l}NL_{s}^{2}d/L)$, highways $O(n_{l}Nd)$; total inference is $\Theta(N)$. At N=8 192N\!=\!8\,192, the subsystem profile in Table 2 shows that the two attention stacks account for $59\%$ of device time, while encoder + decoder remain secondary. Kernel-level profiling in Table 3 then explains why: CUTLASS Tensor-Op GEMMs ($31\%$) and fused attention ($14\%$) dominate, so the workload is primarily compute-bound on Tensor Cores rather than bandwidth-bound on sparse gathers. This directly supports the access-pattern argument in Table 1.
Table 2. Subsystem breakdown of one forward pass at N=8 192N\!=\!8\,192. Attention layers scale with $n_{l}$; everything else is independent of depth.
Subsystem % device time Scales with
Diagonal attention stack $36\%$ $n_{l}$
Off-diagonal attention stack $23\%$ $n_{l}$
Decoder heads $18\%$
Encoder (MLP + GCN) $16\%$
Layout helpers / scatter $4\%$
Table 3. Kernel-level breakdown at N=8 192N\!=\!8\,192. Dominant kernels are compute-bound Tensor-Core GEMMs and fused attention; no SpMV or indexed-gather kernel appears among the top kernels.
Kernel family % device time
CUTLASS Tensor-Op GEMMs (s1688gemm, sm90_xmma, TF32) $31\%$
CUTLASS fused attention (fmha_cutlassF) $14\%$
Elementwise / normalization $17\%$
Copies / reshapes / layout $17\%$
Other $21\%$
5. Preconditioner Application
$M$ is never assembled. The apply consumes the packed factor tensor and produces $M\mathbf{r}$ through three stages on preallocated buffers, with shapes fixed at solve setup by the static partition. (i) A diagonal stage applies $F_{k}$ to each leaf residual $\mathbf{r}_{k}$ in one batched GEMM of shape $(K,L,L)$. (ii) An off-diagonal FMM-style chain makes the data movement explicit:
(2) $\hat{\mathbf{u}}_{k}$ $=\tilde{U}_{k}^{\top}\mathbf{r}_{k},\qquad\hat{\mathbf{v}}_{k}=\tilde{V}_{k}^{\top}\mathbf{r}_{k}$ (restriction)
(3) $\mathbf{s}^{\mathrm{r}}_{m}$ $=\sum_{k\in\mathcal{R}_{m}}\hat{\mathbf{u}}_{k},\qquad\mathbf{s}^{\mathrm{c}}_{m}=\sum_{k\in\mathcal{C}_{m}}\hat{\mathbf{v}}_{k}$ (strip aggregation)
(4) $\mathbf{t}^{\mathrm{c}}_{m}$ $=B_{m}^{\top}\mathbf{s}^{\mathrm{r}}_{m},\qquad\mathbf{t}^{\mathrm{r}}_{m}=B_{m}\mathbf{s}^{\mathrm{c}}_{m}$ (coarse coupling)
(5) $\Delta\mathbf{y}_{k}$ $=\tilde{U}_{k}\,{\textstyle\sum_{m:k\in\mathcal{R}_{m}}\!\mathbf{t}^{\mathrm{r}}_{m}}+\tilde{V}_{k}\,{\textstyle\sum_{m:k\in\mathcal{C}_{m}}\!\mathbf{t}^{\mathrm{c}}_{m}}$ (prolongation).
Stages 1/3/4 are batched GEMMs of shapes $(K,L,L_{s})$ and $(M_{\mathcal{H}},L_{s},L_{s})$ (with $(K,L,L_{s})$ reused in prolongation); stage 2 is two partition-indexed scatter-adds. (iii) The CSR SpMV for $A\mathbf{r}$ is the inner loop’s only sparse operation, and the space-filling-curve ordering keeps gathers nearly banded. Total apply cost is $O(N(L+L_{s}))$ FLOPs. Supplementary LABEL:SX-supp:architecture contains the same equations with full shape progression notes.
Single-graph capture.
Every kernel in the inner loop has fixed launch shape, no host-side allocation, and no data-dependent control flow, so the entire iteration — preconditioner apply, SpMV, SAXPYs, dot-product reductions — records into one CUDA Graph that subsequent iterations replay with a single cudaGraphLaunch. CPU dispatch between kernels disappears, which in a real-time engine usually is the floor on per-frame latency. The graph-capturable property falls out of the design rather than being engineered: IC/DILU triangular solves serialize through data-dependent wavefronts and AMG V-cycles need level-by-level barriers, so neither is single-graph capturable in the same way. Among learned alternatives, neural SPAI (Yang et al., 2025) is graph-capturable but applies $\mathbf{y}\!=\!G^{\top}(G\mathbf{r})$ as two SpMVs against a sparse $G$ whose nonzero pattern follows AA, so its apply remains bandwidth-bound on irregular gathers; in contrast, our factorized approximate-inverse preconditioner apply reduces to dense block GEMMs on the partition and is compute-bound on Tensor Cores.
6. Training
6.1. Cosine Hutchinson Probe Loss
We train the network self-supervised. Given a batch of probe vectors $\mathbf{Z}\in\mathbb{R}^{N\times K_{z}}$ drawn from the spectrally-balanced distribution described below, we compute $A\mathbf{Z}$ via SpMV (treated as a fixed input, with no gradient through AA), apply the preconditioner to obtain $MA\mathbf{Z}$, and minimize the angle between $MA\mathbf{Z}$ and $\mathbf{Z}$ as flattened tensors:
(6) $\mathcal{L}_{\cos}\;=\;1\,-\,\frac{\langle\mathbf{Z},\,MA\mathbf{Z}\rangle_{F}}{\|\mathbf{Z}\|_{F}\,\|MA\mathbf{Z}\|_{F}},$
where $\langle X,Y\rangle_{F}=\operatorname{tr}(X^{\top}Y)$ is the Frobenius inner product and $\|\!\cdot\!\|_{F}$ the Frobenius norm. Equivalently, this is the cosine similarity between $\operatorname{vec}(\mathbf{Z})$ and $\operatorname{vec}(MA\mathbf{Z})$ — a single global angle over all $N\!\cdot\!K_{z}$ entries rather than an average of $K_{z}$ per-probe cosines. The single-denominator form ties all probes together through the same normalization, which we found to give noticeably more stable gradients at small $K_{z}$ than the per-probe mean. A perfect preconditioner gives $\mathcal{L}_{\cos}\!=\!0$; the worst case is $\mathcal{L}_{\cos}\!=\!2$ (anti-aligned).
Why cosine: from vector distance to subspace distance.
The conceptual upgrade $\mathcal{L}_{\cos}$ makes over Frobenius- and SAI-style objectives is to replace a distance between vectors with a distance between subspaces. Frobenius-type losses — including the SAI loss $\bigl\|\tfrac{1}{\|A\|}AM-I\bigr\|_{F}^{2}$ used by (Yang et al., 2025) — penalize the pointwise Euclidean deviation of the specific vector $\tfrac{1}{\|A\|}AM\mathbf{z}$ from the specific vector $\mathbf{z}$. That distance is sensitive to the magnitude of $AM\mathbf{z}$, so it implicitly demands that AMAM act as identity at a particular absolute scale — in the SAI case, at scale $\|A\|$. $\mathcal{L}_{\cos}$, by contrast, sees only the direction of $MA\mathbf{Z}$. Two preconditioners that send a probe to the same one-dimensional subspace incur the same loss, no matter how they scale the vector inside that subspace. Geometrically, $\mathcal{L}_{\cos}$ is a distance on the projective space $\mathbb{P}(\mathbb{R}^{NK_{z}})$ rather than a Euclidean distance on $\mathbb{R}^{NK_{z}}$:
Proposition 0 (Cosine Hutchinson loss is a subspace distance).
Let $M:\mathbb{R}^{N}\!\to\!\mathbb{R}^{N}$ be a linear preconditioner, AA a fixed SPD matrix, and $\mathbf{Z}\in\mathbb{R}^{N\times K_{z}}$ a probe matrix with $\mathbf{Z}\!\neq\!0$ and $MA\mathbf{Z}\!\neq\!0$. Write $\widehat{\mathbf{Z}}\!=\!\operatorname{vec}(\mathbf{Z})$ and $\widehat{\mathbf{Y}}\!=\!\operatorname{vec}(MA\mathbf{Z})$, both elements of $\mathbb{R}^{NK_{z}}$.
  1. (1)
    (Positive-scale invariance.) For every $\alpha\!>\!0$, $\mathcal{L}_{\cos}(\alpha M)\!=\!\mathcal{L}_{\cos}(M)$. The loss therefore descends to a well-defined function on the quotient of preconditioners modulo positive rescaling.
  2. (2)
    (Subspace interpretation.) Let $\theta\in[0,\pi/2]$ be the principal angle between the lines $\mathbb{R}\widehat{\mathbf{Z}}$ and $\mathbb{R}\widehat{\mathbf{Y}}$ in $\mathbb{R}^{NK_{z}}$, and let $\Pi_{\mathbb{R}\mathbf{v}}\!=\!\mathbf{v}\mathbf{v}^{\top}\!/\|\mathbf{v}\|_{2}^{2}$ denote the orthogonal projector onto the line through $\mathbf{v}$. Then
    (7) $\mathcal{L}_{\cos}(M)\;=\;1-\cos\theta,\qquad\tfrac{1}{2}\bigl\|\Pi_{\mathbb{R}\widehat{\mathbf{Z}}}-\Pi_{\mathbb{R}\widehat{\mathbf{Y}}}\bigr\|_{F}^{2}\;=\;1-\cos^{2}\theta,$
    so $\mathcal{L}_{\cos}$ depends $\bigl(\widehat{\mathbf{Z}},\widehat{\mathbf{Y}}\bigr) on$ly through the unsigned angle between the lines they span in $\mathbb{R}^{NK_{z}}$. Both functionals vanish exactly when those two lines coincide.
  3. (3)
    (SAI loss is a vector distance.) The SAI-style loss $\mathcal{L}_{\mathrm{SAI}}(M)\!=\!\bigl\|\,\tfrac{1}{\|A\|}AM\mathbf{Z}-\mathbf{Z}\bigr\|_{F}^{2}$ is not invariant under $M\!\mapsto\!\alpha M$ for $\alpha\!\neq\!1$, and is minimized uniquely (over scalings of $M$) at the choice that places the vector $\tfrac{1}{\|A\|}AM\mathbf{Z}$ as close as possible as a vector in $\mathbb{R}^{NK_{z}}$ to the specific target $\mathbf{Z}$.
Sketch.
(1) Scaling $M$ by $\alpha\!>\!0$ scales the numerator $\langle\mathbf{Z},MA\mathbf{Z}\rangle_{F}$ by $\alpha$ and the denominator $\|\mathbf{Z}\|_{F}\,\|MA\mathbf{Z}\|_{F}$ by the same $\alpha$, leaving the cosine unchanged. (2) The first equality is by definition of $\theta$; the second follows from the rank-one projector identity $\tfrac{1}{2}\|\Pi_{\mathbb{R}\mathbf{u}}-\Pi_{\mathbb{R}\mathbf{v}}\|_{F}^{2}=1-\langle\mathbf{u},\mathbf{v}\rangle^{2}/(\|\mathbf{u}\|^{2}\|\mathbf{v}\|^{2})=1-\cos^{2}\theta$. Both $1\!-\!\cos\theta$ and $1\!-\!\cos^{2}\theta$ are valid notions of squared chordal distance on $\mathbb{P}(\mathbb{R}^{NK_{z}})$ near $\theta\!=\!0$; we use $1-\cos\theta$ in (6) for better-conditioned gradients near the minimum. (3) $\mathcal{L}_{\mathrm{SAI}}$ has the form $\|\beta M\mathbf{u}-\mathbf{u}\|^{2}$ for fixed $\beta\!=\!1/\|A\|$, strictly convex in $\beta M$, hence not rescaling-invariant. ∎
The proposition matches the PCG geometry one-to-one. PCG’s convergence depends only on the relative spread of the eigenvalues of MAMA, not on their absolute location (Sec. 3); accordingly, the correct space to optimize $M$ over is the projective space $M/\mathbb{R}_{>0}$, and the natural loss on that space is a distance between the subspaces $\mathbb{R}\widehat{\mathbf{Z}}$ and $\mathbb{R}\widehat{\mathbf{Y}}$ they span, exactly what $\mathcal{L}_{\cos}$ provides. Frobenius- and SAI-style losses live on the wrong space — they pin down the absolute scale of MAMA even though PCG does not care — and as a side effect implicitly demand that the eigenvalues of MAMA cluster near a chosen value ($\|A\|$ in the SAI case), wasting capacity on a constraint with no algorithmic payoff. By dropping that constraint, $\mathcal{L}_{\cos}$ frees the network to cluster the spectrum wherever the current preconditioner makes angular alignment easiest — a behavior we observe directly in §7.3 (Figs. 42) and ablate against SAI in §7.3.
6.2. Spectrally Biased Probe Vectors
An isotropic Gaussian probe places equal expected energy on every eigenmode of MAMA (Sec. 3), so the resulting gradient signal is also white in the probe’s eigenbasis. The blocks of our preconditioner, however, are not all at the same spatial scale: leaf-diagonal blocks resolve fine, high-frequency structure over $L$ adjacent nodes, while an off-diagonal tile of size SLSL resolves much lower-frequency, larger-scale structure. A spectrally white probe distribution therefore distributes the gradient signal unevenly across these block scales — the high-frequency components, which the fine-scale (diagonal) blocks are tuned to, dominate the signal; the coarse-scale (off-diagonal, large-$S$) blocks receive proportionally weaker gradients. The downstream effect is that blocks at different scales saturate at different times during training, with the coarse blocks plateauing late and limiting overall convergence.
We rebalance the gradient signal across block scales by shifting probe energy toward lower spatial frequencies. A small number of damped-Jacobi smoothing steps acts as a spectral low-pass on the probe:
(8) $\mathbf{z}^{(t+1)}=\mathbf{z}^{(t)}\,-\,\omega\,D^{-1}A\,\mathbf{z}^{(t)},\qquad D=\operatorname{diag}(A),$
with $\omega\!=\!0.6$ and two steps in every reported run. The high-frequency components of the probe are damped more strongly than the low-frequency components, redistributing probe energy toward the eigenmodes the coarse blocks are responsible for. The result is more even gradient magnitudes across block scales and substantially more synchronous training of fine and coarse blocks. Probes are detached, so gradients do not flow back through the smoothing.
6.3. Training Setup
We train with AdamW under a reduce-on-plateau schedule and global gradient clipping. Training contexts (graph, AA, masks, padded sizes, smoothed probes) are precomputed once and cached on disk; at each step a mini-batch is drawn at random and padded to a common node count. We set the number of probe vectors to $K_{z}=\max(64,\lceil\sqrt{N}\rceil)$, balancing gradient noise against compute. The model is compiled with torch.compile. Preconditioner weights and the apply path use float32; only the PCG scalar accumulators (dot products, residual norms, step sizes) are computed in float64, which we found is sufficient to prevent residual drift on stiff systems without paying for full mixed-precision GEMMs.
7. Experiments
7.1. Setup
All GPU experiments run on a single NVIDIA H200. Our model uses d=128d\!=\!128, $n_{l}\!=\!3$, L=128L\!=\!128, $p_{\mathrm{off}}\!=\!4$, $n_{\mathrm{gcn}}\!=\!2$, highways on, trained once for $\sim\!15$ min per scale and reused for every test frame at that scale. All reported solve times are to relative residual $\|\mathbf{r}_{k}\|_{2}/\|\mathbf{r}_{0}\|_{2}\!\leq\!10^{-8}$ — two to three orders of magnitude tighter than the $\sim\!10^{-3}$ tolerance typical of graphics-grade pressure projection, chosen so the ranking reflects preconditioner quality, not early termination. PCG timing uses single-graph CUDA Graph replay for every method that admits it (unpreconditioned CG, Jacobi, AMGX SPAI, neural SPAI, ours) and per-kernel launches for IC/AMG-class methods. AMGX runs with vendor defaults (we swept neighboring configurations at N=8 192N\!=\!8\,192 and saw no improvement). The neural SPAI baseline of Yang et al. (Yang et al., 2025) is re-trained per scale on the same multiphase distribution using their SAI loss and applied as a CUDA-Graph-captured pair of SpMVs. Full hardware/software, precision, and dataset details are in Supplementary §LABEL:SX-supp:experimental-setup.
Benchmark.
We instantiate the target regime — stiff, every-frame-different SPD systems with a hard real-time budget — as 2D multiphase pressure-Poisson: the 5-point Laplacian $A_{ii}\!=\!\sum_{j}w_{ij}$, $A_{ij}\!=\!-w_{ij}$ with harmonic-mean face conductances $w_{ij}\!=\!2\rho_{i}\rho_{j}/(\rho_{i}\!+\!\rho_{j})$, on a per-cell density field $\rho$ randomized per frame across three axes: contrast ($\rho_{\text{heavy}}\!\sim\!\mathrm{loguniform}[5,100]$, so $\kappa(A)\!\in\![10^{3},10^{5}]$), barrier topology (1–3 rectangular barriers with gap configurations including closed, creating near-disconnected sub-domains that force long-range coupling), and orientation. We pick 2D because it is a harder setting for local preconditioners (only 4 neighbors per node) and because it covers the pressure-projection workload that dominates graphics-grade specialized simulators (FLIP/PIC, MPM, fractional-step Navier–Stokes). The architecture is not specific to structured grids — it needs only a sparse graph and a loose spatial ordering — and extends directly to 3D (Fig. 1); we evaluate quantitatively in 2D for fair, fully-tuned comparison against classical baselines. We do not target $N\!\gg\!10^{6}$ regular voxel grids where structured multigrid amortizes (Lyu et al., 2026), nor batched offline PDE workloads. Per-frame randomization and discretization details are in Supplementary §LABEL:SX-supp:dataset; representative frame in Supplementary, Fig. LABEL:SX-supp-fig:bench-frame.
7.2. Main Performance
Table 4. Per-frame end-to-end PCG solve time (ms) and PCG iteration count (in parentheses) on the multiphase Poisson benchmark, 2020 frames per scale, relative residual tolerance $10^{-8}$. GPU methods that admit single-graph capture (unpreconditioned CG, Jacobi, AMGX SPAI, neural SPAI, ours) run the inner loop as a single CUDA Graph replay (§5); IC-class GPU methods and AMGX V-cycles launch per kernel. Our Jacobi/CG numbers already include the CUDA-Graph speedup — an upper bound on what graph capture alone buys without our preconditioner. CPU rows are reference only. Bold marks the fastest GPU time at each scale.
Method 1 0241\,024 2 0482\,048 4 0964\,096 8 1928\,192 16 38416\,384
Unprec. CG (GPU) 18.5 (497) 24.7 (650) 43.6 (1 153) 77.6 (1 765) 89.2 (2 103)
Jacobi (GPU) 12.7 (325) 16.1 (429) 31.7 (839) 39.5 (968) 65.7 (1 543)
AMGX SPAI (GPU) 53.2 (1) 82.0 (1) 76.0 (1) 134.5 (2) 188.3 (3)
IC / DILU (GPU) 139.5 (11) 208.4 (15) 312.8 (22) 503.0 (30) 579.7 (40)
Neural SPAI (GPU) (Yang et al., 2025) 18.4 (118) 26.0 (167) 38.3 (246) 48.1 (338) 70.9 (496)
Ours (GPU) 7.0 (47) 8.8 (66) 9.2 (80) 17.9 (168) 47.6 (394)
CPU reference (not in competition with GPU rows)
IC (CPU) 128.9 (59) 51.6 (96) 159.9 (113) 1 478.9 (170) 13 968.4 (254)
AMG (CPU) 22.2 (5) 24.8 (9) 24.5 (5) 109.8 (7) 674.2 (10)
Neural SPAI (CPU) (Yang et al., 2025) 6.7 (118) 10.9 (167) 25.6 (246) 62.3 (338) 202.4 (496)
Main result.
[…]
Fig. 6 traces how convergence actually looks on a challenging frame — a closed cross-shaped barrier with stiff density contrast chosen to maximize long-range coupling — across unpreconditioned CG, Jacobi, IC, AMG, and ours, with the right-hand side supported only on thin density interfaces. By k=1k\!=\!1, Jacobi and IC have damped the residual only locally around $\mathrm{supp}(b)$, while AMG and ours have already attenuated it across the whole domain — the visual signature of multiscale transport that the highway buffers implement (§4.3). AMG matches the spread but pays in V-cycle synchronization per iteration; ours runs the inner loop as a single CUDA-Graph capture. Iteration counts on this frame ($288/1097/334/19/384$ for unprec/Jacobi/IC/AMG/ours) span nearly two orders of magnitude, but per-iteration cost reverses the ranking for AMG and IC. Standard deviations across the 20 test frames track condition number rather than NN: Jacobi and unprec. CG show $\sigma/\mu\!\approx\!50$$100\%$ at $N\!\geq\!8\,192$ across the $[5,100]$ contrast range, while ours stays at $\leq\!21\%$ even at N=16 384N\!=\!16\,384.
7.3. Training dynamics and ablations
Training dynamics and loss.
Fig. 4 tracks $\mathcal{L}_{\cos}$, the SAI loss on the same checkpoints, and PCG iteration count across training. The three curves move together until step $\sim\!8\,000$, after which the SAI loss rises from $\sim\!10^{-3}$ to $\sim\!0.5$ while $\mathcal{L}_{\cos}$ keeps falling in lockstep with PCG iterations — direct evidence of Prop. 6.1: the model is moving the eigenvalues of MAMA away from $\|A\|$ in pointwise terms while tightening the relative cluster wherever angular alignment is easiest. Fig. 2 confirms this on the spectrum at N=1 024N\!=\!1\,024: SAI delivers a $16\times$ $\kappa$ reduction with the cluster anchored near $\|A\|$, the same architecture trained with $\mathcal{L}_{\cos}$ delivers $68\times$ with the cluster wherever it pleased — a $4.3\times$ gap attributable to the loss alone. Total wall-clock training is $16.2$ min on a single H200 (24 30024\,300 steps); the model overtakes GPU Jacobi by step $\sim\!2\,000$ ($\sim\!1.3$ min) and drops below $200$ PCG iterations by step $\sim\!12\,000$ (Fig. 7 corroborates the link from probe-space alignment to spectral clustering).
Architecture ablations.
Width $d$, depth $n_{l}$, and highway connections each move total solve time non-trivially. Default (d=128d\!=\!128, $n_{l}\!=\!3$, hw on, $\sim\!2.6$M parameters): (i) shrinking depth $n_{l}\!=\!3\!\to\!1$ cuts inference $2.6\times$ but doubles total solve time because iterations rise $2.5\times$ (additional layers are needed to compose information routed through the highway tokens); (ii) removing highways raises PCG iterations $2.3\times$ at N=2 048N\!=\!2\,048 and the penalty grows with NN; (iii) d=64d\!=\!64 is $28\%$ slower overall, d=256d\!=\!256 is competitive at small NN but raises mean PCG iterations from $105$ to $191$ when averaged over $N\!\in\!\{2048,4096,8192\}$. Table 5 summarizes the core sweep results; full rows remain in Supplementary, Table LABEL:SX-supp-tab:ablation.
Table 5. Core architecture ablations (subset of Supplementary Table LABEL:SX-supp-tab:ablation).
Group Configuration Infer. (ms) Iters Total (ms)
Width (avg. $N\!\in\!\{2048,4096,8192\}$) $d{=}64,\;n_{l}{=}3$, hw 3.0 147 15.4
Width (avg. $N\!\in\!\{2048,4096,8192\}$) $d{=}128,\;n_{l}{=}3$, hw 3.1 105 12.0
Width (avg. $N\!\in\!\{2048,4096,8192\}$) $d{=}256,\;n_{l}{=}3$, hw 3.2 191 19.3
Depth (N=8 192N\!=\!8\,192) $d{=}128,\;n_{l}{=}1$, hw 1.3 421 36.9
Depth (N=8 192N\!=\!8\,192) $d{=}128,\;n_{l}{=}3$, hw 3.4 168 17.9
Highways (N=2 048N\!=\!2\,048) $d{=}128,\;n_{l}{=}3$, hw 3.3 66 8.8
Highways (N=2 048N\!=\!2\,048) $d{=}128,\;n_{l}{=}3$, no-hw 2.2 149 13.8
7.4. Generalization and what the loss buys
We probe within-family deployment robustness at N=4 096N\!=\!4\,096 across three shifts — topology (closed barriers withheld from training), contrast (train on $\rho_{\text{heavy}}\!\in\![5,25]$, evaluate on $(25,100]$), and their composition — and compare three systems on identical eval sets: ours, the same architecture trained instead with the SAI loss of (Yang et al., 2025), and neural SPAI (Yang et al., 2025) trained/evaluated under the same split protocol (Table 6). The same-architecture row is a clean loss ablation; the neural SPAI row is a matched-split learned baseline rather than a full-train upper bound.
Table 6. Generalization at N=4 096N\!=\!4\,096 across the four eval distributions, in PCG iterations and speedup vs. Jacobi (mean over 2020 frames, $\mathrm{rtol}\!=\!10^{-8}$). “Jacobi ms” is the baseline denominator used to compute speedup. Our model and the SAI-loss ablation are trained on the complement of each eval cell (held-out OOD setup); neural SPAI (Yang et al., 2025) is trained with the same split protocol (train on the corresponding restricted distribution for each eval row).
Jacobi Ours (cosine loss) Ours arch + SAI loss (Yang et al., 2025) Neural SPAI (Yang et al., 2025)
Eval distribution ms iters speedup iters speedup iters speedup
Full / in-distribution 29.4 82 $3.4\times$ 405 $0.9\times$ 236 $0.9\times$
Closed barriers only 24.7 68 $3.3\times$ 208 $1.3\times$ 258 $0.7\times$
High contrast 24.4 142 $1.9\times$ 175 $1.5\times$ 264 $0.7\times$
Closed ++ high contrast 25.1 147 $1.9\times$ 174 $1.6\times$ 396 $0.5\times$
[…]
8. Discussion and Future Work
The recipe — a weak-admissibility $\mathcal{H}$-matrix prior, a scale-invariant cosine-Hutchinson objective, and a single-graph apply path — is most useful where it targets: stiff, every-frame-different SPD systems with a hard real-time budget. The architecture itself depends on nothing fluid-specific, only a loose spatial ordering of degrees of freedom, and extends directly to 3D (Fig. 1). We expect the largest gains to persist on other SPD families with (i) geometric locality, (ii) frame-to-frame coefficient changes, and (iii) hard real-time budgets (implicit viscosity/diffusion, soft-body and contact dynamics) and smaller gains where a single matrix is reused long enough for heavy classical setup to amortize. Two known limitations are worth flagging: (a) training pre-sizes the partition to a maximum NN, so pushing past it currently requires retraining — a dynamic-partition variant (constant leaf count at $O(N\log N)$ rather than constant leaf size at $\Theta(N)$, §4.2) removes this ceiling at the cost of one extra pooling pass per layer; (b) the $\mathcal{H}$-matrix prior assumes some spatial locality of AA under its indexing, and degrades on operators without a natural spatial coordinate (power-grid Laplacians, social-network matrices) or where the far-field rank does not decay with separation.